# Deep FlexQP: Accelerated Nonlinear Programming via Deep Unfolding

### v1

@@ H.2.1 Shield-MPPI
The method by yin2023shield approximately solves a nonlinear optimization at every MPC step to generate safe controls given a trajectory from a high-level planner such as a model predictive path integral (MPPI) controller. While the main motivation behind this approach is that it is computationally fast, unfortunately, there are a few flaws in that the method has no real guarantees of safety and that the MPPI trajectory is only used to warm-start this second optimization. The main bottleneck preventing us from solving a more complex optimization in real-time is the solver speed. Therefore, this is an application where accelerating optimizers using deep unfolding can shine.
H.2.2 Randomized Problem Scenarios
@@
We use the same Dubins vehicle model as in Section H.1.1. 100 random scenarios are generated by first sampling a random initial and target state uniformly from $\mathcal{U}(-\bar{x},\bar{x})$, where $\bar{x}=(5,5,\pi)$. An obstacle is randomly sampled so its position falls between the initial and target state with a random radius $r$ depending on the distance between the initial and target state: $r\sim\mathcal{U}(0.01,2)*\max(|p_{x}^{\text{target}}-p_{x}^{\text{init}}|,|p_{y}^{\text{target}}-p_{y}^{\text{init}}|)$. The controls are constrained in the ranges $v\in[-10,10]$ and $\omega\in[-5,5]$, enforced by clamping for Shield-MPPI and through the constraints of Equation 63 for our SQP-based method. The reference trajectory at every MPC step is given by running an MPPI controller that samples 10000 trajectories with a look-ahead horizon of 50 timesteps; with a dynamics discretization of $\Delta t=0.05$, this corresponds to a planning horizon of 2.5 seconds ahead. The system experiences zero-mean Gaussian disturbances in its state at every MPC step with standard deviation $(0.05,0.05,0.01)$. Shield-MPPI is allowed to run up to 5 Gauss-Newton iterations per MPC step, while our SQP safety filter is allowed to run up to 5 SQP iterations per MPC step. These thresholds were determined by estimating the max number of iterations that would still allow for real-time control of the system. Collisions in Figure 7 (right) are counted if the state violates the CBF constraint (i.e., intersects the obstacle). Successes are counted if the vehicle reaches within a 0.1 radius of the target state. Figure 10 shows the problem setup and sample trajectories of our approach compared with Shield-MPPI.
@@
Appendix I Performance Comparisons — Small- to Medium-Scale QPs
[…]
@@ Appendix I Performance Comparisons — Small- to Medium-Scale QPs
It is important to note that these results hold only when the direct method is used to solve the linear system. When using an indirect method such as the conjugate gradient (CG) method, converging in fewer iterations is actually a major benefit as each iteration across all the optimizers have roughly the same computational complexity (see Appendix B for discussion and Appendix J for results with the indirect method).
@@
(a) Huber Fitting
(b) Random Linear OCPs
(c) Double Integrator
(d) Oscillating Masses
@@
Appendix J Performance Comparisons — Large-Scale QPs
[…]
@@
Interestingly, the learned optimizers with the data-driven rules for the penalty parameters required many more CG iterations to solve their respective linear systems compared to the traditional optimizers. There seems to be a tradeoff between number of ADMM iterations required to converge vs. the condition number of the coefficient matrices generated at every ADMM iteration. For example, even though Deep FlexQP converges in relatively few iterations, each iteration required an order of magnitude more CG iterations compared to OSQP to solve the linear systems to a similar precision. Using preconditioning with the learned optimizers would likely lead to even greater performance gains over the traditional optimizers.
@@
Appendix K Sample Parameter Prediction Plots
In this section, we compare the $\alpha$, $\rho_{I}$, and $\rho_{E}$ predictions across the traditional and learned optimizers on a few representative optimization problems from Section 5.1. The parameters from the first 20 iterations of the different optimizers are shown in Figure 14. Since Deep OSQP — RLQP Parameterization, Deep OSQP — Improved, and Deep FlexQP output vectors of penalty parameters $\rho_{I}$ and $\rho_{E}$, we plot the mean prediction. Note that OSQP, FlexQP, and Deep OSQP — RLQP Parameterization use a fixed $\alpha$ of $1.6$, $1.6$, and $1.0$, respectively. We also report the infinity norm of the QP residuals to get a better idea of when during the optimization a prediction is being made.
@@
Appendix L Supplementary Results for Generalization Bounds
Figure 15: Generalization bounds for Deep FlexQP trained on 500 oscillating masses QPs. Following Figure 4, this is another case where the generalization bound using the loss Equation 13 is uninformative.
We provide an overview of the generalization bounds training procedure described by saravanos2025deep, which in turn is adapted from the one described by majumdar2021pac. Let ${\bm{s}}=(P,q,G,h,A,b,x^{*},y_{I}^{*},y_{E}^{*})\sim{\mathcal{D}}$ denote a single sample from data distribution $\mathcal{D}$} and let ${\mathcal{S}}=\{{\bm{s}}_{i}\}_{i=1}^{N}$ be a dataset of $N$ samples. Let $\ell({\bm{s}},\theta)\in[0,1]$ be a bounded loss for hypothesis $\theta\sim{\mathcal{P}}$. The true expected loss is defined as
@@
We rely on the following PAC-Bayes bounds that hold with probability $1-\delta$ (majumdar2021pac, Corollary 1):
@@
where $\mathbb{D}^{-1}(p\lvert\rvert c)=\sup\{q\in[0,1]|\mathbb{D}_{KL}({\mathcal{B}}(p)\lvert\rvert{\mathcal{B}}(q)\leq c\}$ is the inverse KL-divergence for Bernoulli random variables ${\mathcal{B}}(p)$ and ${\mathcal{B}}(q)$. The first bound is tighter and is therefore useful for computing the generalization bounds as a numerical certificate of performance, while the second bound has the nice interpretation of a training loss plus a regularization term that depends on the size of the training set and penalizes being off from the prior ${\mathcal{P}}_{0}$. During training, we minimize the second bound using either the loss Equation 13 or Equation 14.
During the evaluation of the tighter generalization bound (after training is complete), since it is difficult to compute the expectation over $\theta\sim{\mathcal{P}}$ in Equation 65, we instead estimate the sample loss using a large number of samples $\{\theta_{j}\}_{j=1}^{M}$ from $\mathcal{P}$$^{*}$:
@@
Combining these bounds results in a final version of the PAC-Bayes bound that holds with probability $1-\delta-\delta^{\prime}$ (majumdar2021pac):
@@
The prior ${\mathcal{P}}_{0}$ for all our models is a stochastic Deep FlexQP trained for 500 epochs on 500 QPs generated from the Random QPs with Equalities problem class using the supervised learning setup from Section 4. We train Deep FlexQP for the generalization bound loss using either Equation 13 or Equation 14 with $\delta=0.009$. We fix a training set for the generalization bounds using 500 problems from the class of interest and train the model for 1000 epochs. We evaluate Equation 69 using $M=10000$ model samples and $\delta^{\prime}=0.001$. Our bounds therefore hold with 99% probability. We report results comparing Equation 13 vs. Equation 14 for LASSO in Figure 4 and for Oscillating Masses in Figure 15. The test loss is computed over 1000 new samples from the target problem class. Overall, the loss Equation 13 results in a less informative bound as it is above all the optimizers, even though the performance in practice can be much better.
@@
Finally, we train Deep FlexQP on 500 QP subproblems generated from a nonlinear predictive safety filter task described in Appendix H using the same procedure to minimize the generalization bound through the loss defined in Equation 14. We use this Deep FlexQP as the QP solver for the predictive safety filter SQP approach described in Section H.2. This provides a numerical certificate on the performance that would not hold for a traditional optimizer.
Appendix M LSTM vs. MLP Policy Comparison
This section presents an ablation analysis on the use of LSTMs to parameterize the policies in both Deep OSQP and Deep FlexQP. The use of LSTMs further leverages the analogy between deep-unfolded optimizers and RNNs, as discussed in monga2021algorithm. Our hypothesis is that the RNN hidden state can encode a compressed history or context from the past optimization variables and residuals, thereby leading to a better prediction of the algorithm parameters to apply at the current iteration. Using an MLP only provides access to information from the current iterate, which could lead to a myopic choice of parameters.
Our results show that LSTMs enhance performance on several problem classes (Figure 18). LSTMs appear to help the most on problems where the active constraints might switch suddenly from one iteration to the next. These types of problems include the machine learning ones, such as SVM, LASSO, and Huber fitting, as well as some of the control problems, like the oscillating masses.
Figure 17: Comparison of MLP vs. LSTM policies. Performance is compared over 1000 test problems.
@@
Appendix N Ablation Analysis on the Use of Lagrange Multipliers in the Supervised Loss
In Figure 20, we compare the performance of Deep FlexQP on different problem classes using the optimality gap loss from Equation 11 and our proposed loss Equation 12. It is evident that our loss including the Lagrange multipliers outperforms the optimality gap loss in all cases, except for the oscillating masses problem class. This could be simply due to the fact that the performance is already approacing 1e10110-10 at 20 iterations, and so small numerical differences play a bigger role. The overall increase in performance using the new loss can be explained by a stronger gradient signal given to Deep FlexQP to learn policies that ensure $\mu_{I}\geq\left\lVert y_{I}^{*}\right\rVert_{\infty}$ and $\mu_{E}\geq\left\lVert y_{E}^{*}\right\rVert_{\infty}$.

### v2

@@ I.2.1 Shield-MPPI
The method by Yin et al. (2023) approximately solves a nonlinear optimization at every MPC step to generate safe controls given a trajectory from a high-level planner such as a model predictive path integral (MPPI) controller. While the main motivation behind this approach is that it is computationally fast, unfortunately, there are a few flaws with the method because it has no real guarantees of safety and the MPPI trajectory is only used to warm-start this second optimization. The main bottleneck preventing us from solving a more complex optimization in real-time is the solver speed. Therefore, this is an application where accelerating optimizers using deep unfolding can shine.
I.2.2 Randomized Problem Scenarios
@@
We use the same Dubins vehicle model as in Section I.1.1. 100 random scenarios are generated by first sampling a random initial and target state uniformly from $\mathcal{U}((-5,-5,-\pi),(5,5,\pi))$. An obstacle is randomly sampled so its position falls between the initial and target state with a random radius $r$ depending on the distance between the initial and target state: $r\sim\mathcal{U}(0.01,2)*\max(|p_{x}^{\text{target}}-p_{x}^{\text{init}}|,|p_{y}^{\text{target}}-p_{y}^{\text{init}}|)$. The controls are constrained in the ranges $v\in[-10,10]$ and $\omega\in[-5,5]$, enforced by clamping for Shield-MPPI and through the constraints of Eq. 56 for our SQP-based method. The reference trajectory at every MPC step is given by running an MPPI controller that samples 10000 trajectories with a look-ahead horizon of 50 timesteps; with a dynamics discretization of $\Delta t=0.05$, this corresponds to a planning horizon of 2.5 seconds ahead. The system experiences zero-mean Gaussian disturbances in its state at every MPC step with standard deviation $(0.05,0.05,0.01)$. Shield-MPPI is allowed to run up to 5 Gauss-Newton iterations per MPC step, while our SQP safety filter is allowed to run up to 5 SQP iterations per MPC step. These thresholds were determined by estimating the max number of iterations that would still allow for real-time control of the system. Collisions in Fig. 7 (right) are counted if the state violates the CBF constraint (i.e., intersects the obstacle). Successes are counted if the vehicle reaches within a 0.1 radius of the target state. Fig. 10 shows the problem setup and sample trajectories of our approach compared with Shield-MPPI.
@@
Appendix J Performance Comparisons — Small- to Medium-Scale QPs
The vanilla and learned optimizers from Section 5.1 are benchmarked on 1000 test QPs from each problem class with the results summarized in Fig. 11. The best version of OSQP is found using a hyperparameter search over the following configurations: fixed parameters for all iterations, adaptive penalty parameters using the OSQP rule, or adaptive penalty parameters using the ADMM rule. Similarly, the best version of FlexQP is found using a hyperparameter search over the following configurations: fixed parameters for all iterations, adaptive penalty parameters using an OSQP-like rule, or adaptive penalty parameters using the ADMM rule. Problems are considered solved when the infinity norm of the QP residuals reaches below an absolute tolerance of $\varepsilon=10^{-3}$. Optimizers are run with no limit on the number of iterations until a timeout of 1 second (1000 ms) is reached. Timings are compared using the normalized shifted geometric mean, which is the factor at which a specific solver is slower than the fastest one (Mittelmann, 2010). We also compare the average number of iterations required to converge as well as the number of coefficient matrix factorizations required to converge to get a sense of where the optimizers are spending the most time. All methods use the direct method to solve their respective linear systems (i.e., equality-constrained QPs) at every iteration, and the factorization from the previous iteration is reused if the parameters have not changed by more than a factor of 5x following the heuristic used by OSQP (Stellato et al., 2020).
@@
It is important to note that these results hold only when the direct method is used to solve the linear system. When using an indirect method such as the conjugate gradient (CG) method, converging in fewer iterations is actually a major benefit as each iteration across all the optimizers have roughly the same computational complexity (see Appendix C for discussion and Appendix K for results with the indirect method).
@@
(a) Support Vector Machines
(b) LASSO
(c) Huber Fitting
(d) Random Linear OCPs
(e) Double Integrator
(f) Oscillating Masses
@@
Appendix K Performance Comparisons — Large-Scale QPs
In this section, we report the full details of the large-scale QP experiments introduced in Section 5.2. Due to memory limitations, each optimizer is unfolded for 10 iterations and trained with a batch size of 1. Optimizers are trained for 5 epochs on 100 training problems and evaluated on 100 new test problems, with results presented in Fig. 13. Convergence is declared when the infinity norm of the QP residuals falls below an absolute tolerance of $\varepsilon=10^{-3}$. Each optimizer is run until convergence or until a timeout of 10 minutes is reached. Similar to Appendix J, we report the normalized shifted geometric mean to better highlight the relative performance of each optimizer (wall clock times are provided in Fig. 6). We also report the percent of problems solved, the average number of iterations taken to converge, and the average number of CG iterations necessary to solve the linear system at each ADMM iteration. For the optimizers that failed to converge on any problems, we report the average final QP residual infinity norms and average number of iterations run in Tables 2 and 3. This gives a rough idea of how far away the optimizers were from converging when they were timed out.
@@
Interestingly, the learned optimizers with the data-driven rules for the penalty parameters required many more CG iterations to solve their respective linear systems compared to the traditional optimizers. There seems to be a tradeoff between number of ADMM iterations required to converge vs. the condition number of the coefficient matrices generated at every ADMM iteration. For example, even though Deep FlexQP converges in relatively few iterations, each iteration requires an order of magnitude more CG iterations compared to OSQP to solve the linear systems to a similar precision. Using preconditioning with the learned optimizers would likely lead to even greater performance gains over the traditional optimizers.
Appendix L Sample Parameter Prediction Plots
@@
In this section, we compare the $\alpha$, $\rho_{I}$, and $\rho_{E}$ predictions across the traditional and learned optimizers on a few representative optimization problems from Section 5.1. The parameters from the first 20 iterations of the different optimizers are shown in Fig. 14. Since Deep OSQP — RLQP Parameterization, Deep OSQP — Improved, and Deep FlexQP output vectors of penalty parameters $\rho_{I}$ and $\rho_{E}$, we plot the mean prediction. Note that OSQP, FlexQP, and Deep OSQP — RLQP Parameterization use a fixed $\alpha$ of $1.6$, $1.6$, and $1.0$, respectively. We also report the infinity norm of the QP residuals to get a better idea of when during the optimization a prediction is being made.
Appendix M Supplementary Results for Generalization Bounds
Figure 15: Generalization bounds for Deep FlexQP trained on 500 oscillating masses QPs. Following Fig. 4, this is another case where the generalization bound using the loss Eq. 13 is uninformative.
We provide an overview of the generalization bounds training procedure described by Saravanos et al. (2025), which in turn is adapted from the one described by Majumdar et al. (2021). Let ${\bm{s}}=(P,q,G,h,A,b,x^{*},y_{I}^{*},y_{E}^{*})\sim{\mathcal{D}}$ denote a single sample from data distribution $\mathcal{D}$} and let ${\mathcal{S}}=\{{\bm{s}}_{i}\}_{i=1}^{N}$ be a dataset of $N$ samples. Let $\ell({\bm{s}},\theta)\in[0,1]$ be a bounded loss for hypothesis $\theta\sim{\mathcal{P}}$. The true expected loss is defined as
@@
We rely on the following PAC-Bayes bounds that hold with probability $1-\delta$ (Majumdar et al., 2021, Corollary 1):
@@
where $\mathbb{D}^{-1}(p\lvert\rvert c)=\sup\{q\in[0,1]|\mathbb{D}_{KL}({\mathcal{B}}(p)\lvert\rvert{\mathcal{B}}(q)\leq c\}$ is the inverse KL-divergence for Bernoulli random variables ${\mathcal{B}}(p)$ and ${\mathcal{B}}(q)$. The first bound is tighter and is therefore useful for computing the generalization bounds as a numerical certificate of performance, while the second bound has the nice interpretation of a training loss plus a regularization term that depends on the size of the training set and penalizes being off from the prior ${\mathcal{P}}_{0}$. During training, we minimize the second bound using either the loss Eq. 13 or Eq. 14.
During the evaluation of the tighter generalization bound (after training is complete), since it is difficult to compute the expectation over $\theta\sim{\mathcal{P}}$ in Eq. 58, we instead estimate the sample loss using a large number of samples $\{\theta_{j}\}_{j=1}^{M}$ from $\mathcal{P}$$^{*}$:
@@
Combining these bounds results in a final version of the PAC-Bayes bound that holds with probability $1-\delta-\delta^{\prime}$ (Majumdar et al., 2021):
@@
The prior ${\mathcal{P}}_{0}$ for all our models is a stochastic Deep FlexQP trained for 500 epochs on 500 QPs generated from the Random QPs with Equalities problem class using the supervised learning setup from Section 4. We train Deep FlexQP for the generalization bound loss using either Eq. 13 or Eq. 14 with $\delta=0.009$. We fix a training set for the generalization bounds using 500 problems from the class of interest and train the model for 1000 epochs. We evaluate Eq. 62 using $M=10000$ model samples and $\delta^{\prime}=0.001$. Our bounds therefore hold with 99% probability. We report results comparing Eq. 13 vs. Eq. 14 for LASSO in Fig. 4 and for Oscillating Masses in Fig. 15. The test loss is computed over 1000 new samples from the target problem class. Overall, the loss Eq. 13 results in a less informative bound as it is above all the optimizers, even though the performance in practice can be much better.
Finally, we train Deep FlexQP on 500 QP subproblems generated from a nonlinear predictive safety filter task described in Appendix I using the same procedure to minimize the generalization bound through the loss defined in Eq. 14. We use this Deep FlexQP as the QP solver for the predictive safety filter SQP approach described in Section I.2. This provides a numerical certificate on the performance that would not hold for a traditional optimizer.
@@
Appendix N LSTM vs. MLP Policy Comparison
This section presents an ablation analysis on the use of LSTMs to parameterize the policies in both Deep OSQP and Deep FlexQP. The use of LSTMs further leverages the analogy between deep-unfolded optimizers and RNNs, as discussed in Monga et al. (2021). Our hypothesis is that the RNN hidden state can encode a compressed history or context from the past optimization variables and residuals, thereby leading to a better prediction of the algorithm parameters to apply at the current iteration. Using an MLP only provides access to information from the current iterate, which could lead to a myopic choice of parameters.
Our results show that LSTMs enhance performance on several problem classes (Fig. 17). LSTMs appear to help the most on problems where the active constraints might switch suddenly from one iteration to the next. These types of problems include the machine learning ones, such as SVM, LASSO, and Huber fitting, as well as some of the control problems, like the oscillating masses.
@@
Appendix O Ablation Analysis on the Use of Lagrange Multipliers in the Supervised Loss
In Fig. 19, we compare the performance of Deep FlexQP on different problem classes using the optimality gap loss from Eq. 11 and our proposed loss Eq. 12. It is evident that our loss including the Lagrange multipliers outperforms the optimality gap loss in all cases, except for the oscillating masses problem class. This could be simply due to the fact that the performance is already approaching $10^{-10}$ at 20 iterations, and so small numerical differences play a bigger role. The overall increase in performance using the new loss can be explained by a stronger gradient signal given to Deep FlexQP to learn policies that ensure $\mu_{I}\geq\left\lVert y_{I}^{*}\right\rVert_{\infty}$ and $\mu_{E}\geq\left\lVert y_{E}^{*}\right\rVert_{\infty}$.